# Enhancing the NAO: Extending Capabilities of Legacy Robots for Long-Term Research

### v1

@@ IV-A Hardware
In addition to the NAO’s two head-mounted cameras, we integrated a depth and thermal camera, expanding multimodal vision with synchronized RGB-D and thermal streams. The Hailo-8L accelerator allows real-time inference with custom models on these streams, a capability unavailable on the NAO AI. The ReSpeaker handles audio capture, rather than the NAO v6’s built-in microphones, bypassing NAOqi’s restrictions. NAOqi provides audio only through ALAudioDevice in either 16 kHz or 48 kHz in 170 ms chunks [1], without support for beam steering or noise suppression, limiting custom ASR pipelines.
@@ IV-B Software
We restructured the conversational pipeline by replacing the NAO’s intent-driven dialogue system with real-time, cloud-based transcription and diarization (Deepgram [36]). The Enhanced NAO uses an LLM, rather than fixed templates, for response generation, which enables dynamic, free-flowing conversation. Finally, to produce natural-sounding voices, we use Amazon Polly [37] for text-to-speech (TTS) generation.
@@ IV-C Conversation
All conversations were coded by a primary researcher, with 20% independently coded by a secondary researcher to assess reliability. Inter-rater reliability [42], calculated using a two-way random intraclass correlation coefficient (ICC[2,1]), indicated excellent agreement between coders ($\text{ICC}=0.9897$).
For each of the four conversation attributes (average response time, conversational quality, turn balance, and number of topic shifts), we performed paired t-tests with a Bonferroni correction ($\alpha=0.0125$) to account for multiple comparisons (Fig. 5). We report effect sizes using Cohen’s $d$. Average response time did not differ significantly between the Enhanced NAO ($M=1.94,SD=0.42$) and the NAO AI ($M=2.13,SD=0.30$), $t(14)=1.05,p=0.3105,d=0.53$. Conversational quality was significantly higher for the Enhanced NAO ($M=2.89,SD=0.11$) compared to the NAO AI ($M=2.11,SD=0.37$), t(14)=5.69,p<0.001,d=2.85t(14)=-5.69,p<0.001,d=-2.85. Turn balance, was also significantly higher for the Enhanced NAO ($M=4.00,SD=1.19$) than the NAO AI ($M=0.86,SD=0.21$), t(14)=6.89,p<0.001,d=3.44t(14)=-6.89,p<0.001,d=-3.44, indicating that the Enhanced NAO produced substantially more words relative to the user. Additionally, the Enhanced NAO exhibited fewer topic shifts ($M=0.12,SD=0.35$) than the NAO AI ($M=1.44,SD=1.35$), $t(14)=2.66,p=0.0185,d=1.33$, which approached significance after the Bonferroni Correction.
@@ V Discussion
These significant hardware upgrades directly improved the conversational quality of the Enhanced NAO. Cleaner, lower-latency audio from the ReSpeaker and the LLM-driven conversational pipeline generated more natural, context-aware responses. In contrast, the NAO AI’s rigid intent-based models produced repetitive, unnatural, and disruptive replies, reducing flow. Response times did not differ significantly between conditions, showing that richer interactions can be delivered without added latency, despite additional sensing, processing, and cloud services. Interestingly, turn balance skewed toward the Enhanced NAO, reflecting its ability to maintain a conversation, but also underscoring the need to preserve balance in future designs. Importantly, the Enhanced NAO showed fewer topic shifts, suggesting conversations stayed more coherent and user-driven.
@@ V-A Limitations
A second limitation involved a technical inconsistency in one session, where a participant interacted with the Enhanced NAO using a male TTS voice instead of the female voice presented in all other sessions. While this mismatch may have influenced the system’s performance and the participant’s perception of the robot’s gender, it also revealed the flexibility of our pipeline to support multiple voices. Interestingly, this participant was the only one to prefer the NAO AI over the Enhanced NAO during qualitative feedback, which may have been due to the technical issue or the use of the male voice. Future experiments will ensure identical conditions across participants to eliminate such confounds.
Third, because this study emphasized integration over conversational design, aspects such as response prompt engineering and pretesting were not fully optimized, occasionally producing responses that were overly long or interview-like. These design refinements are straightforward to implement and will be addressed in subsequent studies, which will employ larger participant pools, standardized conditions, and improved dialogue strategies to rigorously evaluate system effectiveness.
Finally, we did not formally evaluate the new multimodal camera system in this pilot. Integration was the focus of this pilot, while assessing perceptual and interaction benefits will be pursued in future studies. These limitations reflect the preliminary nature of the study, but also underscore its success in validating the technical feasibility of our integrated framework and establishing the groundwork for more comprehensive evaluation.
@@ VI Conclusion and future work
Future work will address the remaining technical challenges to further improve interaction quality. In particular, we will migrate our speech-to-text and text-generation pipelines from cloud services to local processing using compact, high-performance models such as Gemma 3n [43] or Qwen 3 [44]. Running these models locally is expected to further reduce latency, strengthen privacy, and increase system autonomy. We will also leverage multimodal visual perception to detect and localize users prior to speech, enabling natural dialogue initiation. In multi-party settings, fused visual and thermal data will support more accurate speaker separation by disambiguating overlapping voices and tracking active participants, further improving responsiveness.

### v2

@@ IV-A Hardware
In addition to the NAO’s two head-mounted cameras, we integrated a depth and thermal camera, expanding multimodal vision with synchronized RGB-D and thermal streams. The Hailo-8L accelerator allows real-time inference with custom models on these streams, a capability unavailable on the NAO AI. The ReSpeaker handles audio capture, rather than the NAO v6’s built-in microphones, bypassing NAOqi’s restrictions. NAOqi provides audio only through ALAudioDevice in either 16 kHz or 48 kHz in 170 ms chunks [aldebaran_nao_2025], without support for beam steering or noise suppression, limiting custom ASR pipelines.
@@ IV-B Software
We restructured the conversational pipeline by replacing the NAO’s intent-driven dialogue system with real-time, cloud-based transcription and diarization (Deepgram [deepgram2025transcription]). The Enhanced NAO uses an LLM, rather than fixed templates, for response generation, which enables dynamic, free-flowing conversation. To produce natural-sounding voices, we use Amazon Polly [AmazonPolly_2025] for text-to-speech (TTS) generation. These software modifications are part of the system’s functional improvements, which directly contribute to the Enhanced NAO’s observed gains in conversational quality and motivated the selection of our evaluation metrics.
@@ IV-C Conversation
As described in Sec. III, all conversations were independently coded by two researchers. Inter-rater reliability [hallgren2012computing], calculated using a two-way random intraclass correlation coefficient (ICC[2,1]), indicated excellent agreement between coders ($\text{ICC}=0.9897$).
For each conversation attribute (average response time, conversational quality, turn balance, and number of topic shifts), we performed a paired t-test with a Bonferroni correction ($\alpha=0.0125$) to account for multiple comparisons (Fig. 6). We report effect sizes using Cohen’s $d$. Average response time did not differ significantly between the Enhanced NAO ($M=1.94,SD=0.42$) and the NAO AI ($M=2.13,SD=0.30$), $t(14)=1.05,p=0.3105,d=0.53$. Conversational quality was significantly higher for the Enhanced NAO ($M=2.89,SD=0.11$) compared to the NAO AI ($M=2.11,SD=0.37$), t(14)=5.69,p<0.001,d=2.85t(14)=-5.69,p<0.001,d=-2.85. Turn balance, was also significantly higher for the Enhanced NAO ($M=4.00,SD=1.19$) than the NAO AI ($M=0.86,SD=0.21$), t(14)=6.89,p<0.001,d=3.44t(14)=-6.89,p<0.001,d=-3.44, indicating that the Enhanced NAO produced substantially more words relative to the user. Additionally, the Enhanced NAO exhibited fewer topic shifts ($M=0.12,SD=0.35$) than the NAO AI ($M=1.44,SD=1.35$), $t(14)=2.66,p=0.0185,d=1.33$, which approached significance after the alpha correction.
@@ V Discussion
These significant hardware upgrades directly improved the conversational quality of the Enhanced NAO. Cleaner, lower-latency audio from the ReSpeaker and the LLM-driven conversational pipeline generated more natural, context-aware responses. In contrast, the NAO AI’s rigid intent-based models produced repetitive, unnatural, and disruptive replies, reducing flow. Response times did not differ significantly between conditions, showing that richer interactions can be delivered without added latency, despite additional sensing, processing, and cloud services. Interestingly, turn balance for the Enhanced NAO exceeded our desired range, yet participants still preferred it. This suggests that turn balance may be less critical to user preference than we initially assumed, or that our acceptable range was too conservative. Nevertheless, avoiding overly long or monologic robot responses is still necessary for maintaining enjoyable, user-centered conversation. Additionally, the Enhanced NAO exhibited fewer topic shifts, suggesting conversations stayed more coherent and user-driven.
While the primary focus of this work was technical enhancement, we also acknowledge that the visible mechanical additions may have influenced user perception. The NAO’s humanoid, friendly appearance supports comfort and approachability, especially among children [shamsuddin2012humanoid, amirova202110, rossi2022using]. Exposing additional hardware elements may make the robot appear more mechanical and potentially reduce approachability. Future adaptations should therefore consider visual and tactile design elements. Prior research on social and affective robots highlights the benefits of softness and warmth for comfort during social-physical interactions [block2019softness, block2017should, block2018emotionally, block2021six, block2023arms, block2021huggiebot]. Inspired by this line of work, future iterations of the Enhanced NAO could incorporate a soft cover or hat over exposed components to preserve the NAO’s approachability while maintaining the improved functionality.
@@ V-A Limitations
A second limitation involved a technical inconsistency in one session, where a participant interacted with the Enhanced NAO using a male TTS voice instead of the female voice presented in all other sessions. This mismatch may have influenced the system’s performance and the participant’s perception of the robot’s gender, and notably, this was the only participant who preferred the NAO AI. The deviation could also have interacted with turn-balance patterns. However, it also revealed the flexibility of our pipeline to support multiple voices. We did not analyze male vs. female voice as an experimental factor in this pilot, but future work should explicitly test how voice characteristics influence conversational quality and user preference. Future experiments will also implement stricter quality control to ensure consistent conditions across participants.
Additionally, participants were not blinded to the experimental condition. To mitigate bias, we used a counter-balanced condition order and asked participants to evaluate the conversation quality rather than the robot itself. Future studies could incorporate blinded or partially blinded setups to minimize potential bias due to visual differences between conditions.
Forth, because this pilot emphasized system integration over conversational design, prompt engineering and pretesting were not fully optimized, occasionally producing overly long or interview-like responses. These refinements are straightforward to implement and will be incorporated in future studies with larger participant pools, standardized conditions, and improved dialogue strategies.
Finally, although the new multimodal camera system was successfully integrated, we did not formally evaluate its perceptual or interaction benefits in this pilot. Integration was the primary goal; follow-up work will characterize the contributions of these sensing modalities in depth. Taken together, these limitations reflect the preliminary nature of the study, but also underscore its success in validating the technical feasibility of our integrated framework and establishing the groundwork for more comprehensive evaluation.
@@ VI Conclusion and future work
Future work will address the remaining technical challenges to further improve interaction quality. In particular, we will migrate our speech-to-text and text-generation pipelines from cloud services to local processing using compact, high-performance models such as Gemma 3n [gemma_3n_2025] or Qwen 3 [qwen3technicalreport]. Running these models locally is expected to further reduce latency, strengthen privacy, and increase system autonomy. We will also leverage multimodal visual perception to detect and localize users prior to speech, enabling natural dialogue initiation. In multi-party settings, fused visual and thermal data will support more accurate speaker separation by disambiguating overlapping voices and tracking active participants, further improving responsiveness.